OpenStudy (anonymous):

WILL GIVE MEDAL IF YOU HELP WITH QUADRATIC FUNCTIONS What are the exact solutions of x2 = 4 − 7x?

OpenStudy (hlambach):

is the x2 x raised to the power of two?

OpenStudy (anonymous):

yes : )

OpenStudy (hlambach):

Remember, quadratic function is in the form 0 = ax^2 + bx + c After you have put the formula in that form, simply use the quadratic formula: \[x = \frac{ -b \pm \sqrt{b ^{2} - 4ac} }{ 2a }\]

OpenStudy (anonymous):

im so confused.. i dont understand these at all... my answer choices are

OpenStudy (anonymous):

x = x equals negative 7 plus or minus the square root of thirty-three all over 2 x = x equals negative 7 plus or minus the square root of sixty-five all over 2 x = x equals 7 plus or minus the square root of sixty-five all over 2 x = x equals 7 plus or minus the square root of thirty-three all over 2

OpenStudy (hlambach):

Don't worry, I'll guide you.

OpenStudy (hlambach):

To get your original equation in quadratic form, simply subtract x^2 from both sides, so you're left with 0 = 1x^2 - 7x + 4 (I rearranged it so it's easier to follow.) Now, plug in the numbers in the equation I gave you above, but not in a calculator.